INSURANCE COMPANY Z REIMBURSES PHARMACIES AT AWP+2%+$4.50 PRESCRIPTION. hOW MUCH WILL THEY PAY FOR 120 TABLETS OF TAGAMET(AWP=$92/100)?

1
Expert's answer
2021-11-07T19:45:13-0500

Solution:

AWP = 92/100 = 0.92

= 0.92 + 2% + 4.50 = 0.92 + 0.02 + 4.50 = $5.44 per tablet

For 120 tablets: $5.44 x 120 = $652.8

They will pay = $653